OpenStudy (anonymous):

How many ways are there to seat 4 people in a 9-passenger automobile? Assume that some times there will be no one in the driver's seat.

OpenStudy (anonymous):

we have 4 person and 9 seat, first one can choose from 9 seats, second 8, third 7, fourth from 6, so there are 9*8*7*6 ways -wekku

OpenStudy (anonymous):

so 3024 ways?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I disagree.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

It would only be that many if we care who is sitting in which seat.

OpenStudy (anonymous):

How many ways are there to seat 4 people in a 6-passenger automobile? Assume that some times there will be no one in the driver's seat. what about this question? same type just different numbers

OpenStudy (anonymous):

yeah. same type deal. I think you are being asked to do combinations again. You have a pool of 9 seats. You will be picking 4 of the 9 seats to fill.

OpenStudy (anonymous):

So how many different combinations of 4 seats can you make from 9 seats?

OpenStudy (anonymous):

ummmm i'm not sure

OpenStudy (anonymous):

Recall our previous notation?

OpenStudy (anonymous):

How big is the pool?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

i'm talking about the six passenger car

OpenStudy (anonymous):

Ok, and how many seats will we pick for the combinations?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

So the answer in nCr notation is?

OpenStudy (anonymous):

30?

OpenStudy (anonymous):

No, write the nCr notation first.

OpenStudy (anonymous):

6!/4!(6!-6)!

OpenStudy (anonymous):

Close!

OpenStudy (anonymous):

\[\frac{6!}{4!(6-4)!}\]

OpenStudy (anonymous):

ahhhhhh! i got it mixed up!

OpenStudy (anonymous):

so 60?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

\[\frac{6!}{4!(6-4)!} = \frac{6!}{4!2!} = \frac{6\times 5}{2!} = \frac{30}{2} = 15\]

OpenStudy (anonymous):

so the answer is 15?

OpenStudy (anonymous):

maybe. Did you check my work?

OpenStudy (anonymous):

it should be 360

OpenStudy (anonymous):

how so?